Show that Z[√-2] is not a UFD.
1
Expert's answer
2020-03-03T16:51:22-0500

Actually, "\mathbb{Z}[\sqrt{-2}]" is UFD.

Proof:

We will prove that "\mathbb{Z}[\sqrt{-2}]" is Euclidean domain. It will be enough because every Euclidean domain is a unique factorization domain.

If there exists a Euclidean norm on "\mathbb{Z}[\sqrt{-2}]" , then the integral domain "\mathbb{Z}[\sqrt{-2}]" is a Euclidean domain. (definition)

So, we will prove, that in the integral domain "\mathbb{Z}[\sqrt{-2}]" , the function "N(\alpha)=\alpha \overline{\alpha}" is a multiplicative Euclidean norm.

(i) "\forall r\in \mathbb{Z}[\sqrt{-2}] /\{0\} : N(r)\geq0"

"N(a+b\sqrt{-2})=(a+b\sqrt{-2})(a-b\sqrt{-2})=a^2+2b^2\geq 0"


(ii) "\forall a,b\in \mathbb{Z}[\sqrt{-2}] , b \cancel{=}0, \ \exists r,q\in \mathbb{Z}[\sqrt{-2}] : a=bq+r \ (r=0" or "N(r)<N(b))"

Let "a+b\sqrt{-2}, \ c+d\sqrt{-2} \in \mathbb{Z}[\sqrt{-2}] , \quad a,b,c,d\in \mathbb{Z}."

"\exists e,f\in \mathbb{Q}: \quad \frac{a+b\sqrt{-2}}{c+d\sqrt{-2}}=\frac{(ac+2bd)+(bc-ad)\sqrt{-2}}{c^2+2d^2}=e+f\sqrt{-2}"

Consider "r,s\in \mathbb{Z}: \ |e-r|\leq1/2, |f-s|\leq1/2"

Then "a+b\sqrt{-2}=(c+d\sqrt{-2})(e+f\sqrt{-2})="

"=(c+d\sqrt{-2})(r+s\sqrt{-2}+(e-r)+(f-s)\sqrt{-2})="

"=(c+d\sqrt{-2})(r+s\sqrt{-2})+(c+d\sqrt{-2})((e-r)+(f-s)\sqrt{-2})"

We need to check that "N(c+d\sqrt{-2})>N((c+d\sqrt{-2})((e-r)+(f-s)\sqrt{-2})" .

"N((c+d\sqrt{-2})((e-r)+(f-s)\sqrt{-2}))= N(c+d\sqrt{-2})N((e-r)+(f-s)\sqrt{-2})\leq N(c+d\sqrt{-2})(1/4+2\times 1/4)=3/4 N(c+d\sqrt{-2})< N(c+d\sqrt{-2})"



Therefore, "N" is Euclidean norm "\Rightarrow" "\mathbb{Z}[\sqrt{-2}]" is Euclidean domain "\Rightarrow" "\mathbb{Z}[\sqrt{-2}]" is UFD.



This material can be found here: https://math.stanford.edu/~ksound/Math152A10/Solution6.pdf